A right triangle has a hypotenuse of length 25 and a leg of length 20. What is the length of the other leg? If necessary, round your answer to two decimal places.
Answer by dabanfield(803)

A right triangle has a hypotenuse of length 25 and a leg of length 20. What is the length of the other leg? If necessary, round your answer to two decimal places.
Let x be the unknown side. Then by the Pythagorean Theorm we have:

25^2 = 20^2 + x*2
625 = 400 + x^2
x^2 = 625 - 400 = 225
x = sqrt(225) = 15